Mr. Rodney's English class is made up of 28 students. He has 6 ESL students, 10 remedial student, and 5 advanced learners. ESL students make up 1/5 of the remedial students and 3/5 of the advanced learners.
1.) What is the probability that a student is ESL and remedial?
2.) What is the probability that a student is ESL and an advanced learner?
3.) What is the probability that a student is remedial and NOT ESL?
Thank you for your help.

Answer by edjones(7311)

r=remedial only, a=advanced only, e=esl only, s=students who are not r, e, or a
8 r
2 re
2 a
3 ae
1 e
12 s
----add
28
.
1) 2/28 = 1/14
2) 3/28
3) 8/28 = 2/7
.
Ed